OpenStudy (anonymous):

absoulute value equations |1\2x+2|=8

OpenStudy (anonymous):

subtract 2 from both sides 1/2x=6 invert the fraction and multiply x= absolute value of 12

OpenStudy (shamil98):

There are two solutions.

OpenStudy (shamil98):

1/2 x + 2 = 8 1/2x = 6 x = 12 1/2x + 2 = -8 1/2x = -10 x = -20